OpenStudy (anonymous):

Which of the following equations is written in slope-intercept form? 2x + 4y = 8 4y = -2x + 8 y = -1/2x + 2 x = -2y + 4

OpenStudy (anonymous):

slope intercept form would be y=mx+c

OpenStudy (anonymous):

so option 3

OpenStudy (hoblos):

y = -1/2x + 2

OpenStudy (anonymous):

y=mx+b not c & the third option(:

OpenStudy (anonymous):

y = -1/2x + 2

OpenStudy (anonymous):

Choose the correct procedure used to graph the equation 3x + 9y = -18 using the intercepts method. Solve the equation for y to get a y-intercept of (0, -2) and a slope of -1/3. Plot the y-intercept, count the slope to get another point and draw the line that passes through the two points. Choose two different values; plug them into the equation to find the corresponding y-values. Plot the two points and draw the line that passes through the two points. Let x equal zero and solve for y to find a y-intercept of (0, -2). Let y equal zero and solve for x to find an x-intercept of (-6, 0). Plot the two points and draw a line passing through the two points. Solve the equation for y to find an intercept of (-2, 0); solve the equation for x to find a y-intercept of (0, -6). Plot the two points and draw a line passing through the two points.